A rectangle has a perimeter of 34 cm and an area of 52 cm. Its length is 5 more than twice its width. Write and solve a system of equations to find the dimensions of the rectangle.
Answer by oberobic(2304)

2(L+W) = 34
L + W = 17
.
L*W = 52
.
L = 2W+5
.
2L + 2W = 34
L - 2W = 5
--------------- add
3L = 39
L = 13
.
L+W = 17
W = 4
.
Check perimeter.
2*(13+4) = 34
Check area.
13*4 = 42
Check relationship between L and W.
L = 2*W + 5 ??
Yes.
.
Answer: Length = 13; Width 4
.
Done.